(1) Computer hosts usually have two addresses, an Internet Protocol (IP) address and an Ethernet Media Access Control (MAC) address. For the benefit of humans, the IP address is normally represented as a dotted decimal notation, such as 192.168.10.2. Each of the decimal octets in the address or a mask can be converted to 8 binary bits. Remember that the computer only understands binary bits. If all 4 octets were converted to binary, how many bits would there be?

(2) Manipulate powers of 2 to determine the number of hosts on a network

  1. a. Binary numbers use two digits, 0 and 1. When you calculate how many hosts can be on a subnetwork, you use powers of two because binary is being used. As an example, we have a subnet mask that leaves six bits in the host portion of the IP address. In this case, the number of hosts on that network is 2 to the 6th power minus 2 (because you need a number to represent the network and a number that can be used to reach all the hosts—the broadcast address). The number 2 is always used because we are working in binary. The number 6 is the number of bits that are used for the host bits.
  2. b. Change the calculator view to Scientific mode. Input the number 2. Select the xy key on the calculator, the key which raises a number to a power. Input the number 6. To compete the operation, click on the = key, press Enter on the keyboard, or press the = key on the keyboard. The number 64 appears in the output. To subtract two, click on the minus (-) key and then the 2 key followed by the = key. The number 62 appears in the output. This means 62 hosts could be utilized.
  3. c. Using the previously described process, determine the number of hosts if the following number of bits are used for host bits.

No. of Bits Used for Hosts

No. of Hosts

5

14

24

10

(3) List one other thing for which you might use the Windows Calculator scientific mode. It does not have to be related to networking.
Answer #1

Answer 1
**********

If all 4 octets were converted to binary, how many bits would there be?

32 bits as each octet contain 8 bits so 8x4=32 bits is the length of IPv4

Answer 2
*********
generic formula for host calculation 2^n-2

2^5= 32-2 = 30
2^14 = 16384 - 2 = 16382
2^24 = 16777216 - 2 = 16777214
2^10 = 1024 - 2 = 1022

Answer 3
*********
could be anything like mod,log, or to square any number

if you have any doubt then please ask me without any hesitation in the comment section below , if you like my answer then please thumbs up for the answer , before giving thumbs down please discuss the question it may possible that we may understand the question different way and we can edit and change the answers if you argue, thanks :)